OpenStudy (anonymous):

(-5/7)(5) = Answer -5/35 -1/7 -3 4/7 -2 6/7

OpenStudy (anonymous):

@evergirl

OpenStudy (anonymous):

@privetek

OpenStudy (anonymous):

so with this one you would set it up like (-5/7) * (5/1) = -25/7. simplify it from there.

OpenStudy (anonymous):

-5/35 @PaulK

OpenStudy (anonymous):

sorry, yo no se.....

OpenStudy (anonymous):

nope you would set it up as in mulitplying the top number to top number and bottom bnumber to bottom number. so it would be (-5 * 5) / (7*1) = -25/7. but you have to simplify it from there it looks like because that doesn't match your answers above in this form. so go from there.

OpenStudy (anonymous):

do you have it now?

OpenStudy (anonymous):

-25/7 divide 25/7  remainder is 4 so... -3 4/7